OpenStudy (pinkkjazzy3):

Write the sum of 6 and -8 as a complex number in standard form.

OpenStudy (perl):

assuming 6 and -8 are real numbers 6 = 6 + 0*i -8 = -8 + 0*i

OpenStudy (pinkkjazzy3):

it would be -8+0i=-8?